# Cash Flow Underwriting with Bank Transaction Data: Advancing MSME Financial Inclusion in Malaysia

### v3

@@ 7 Deployment and Operational Framework
The sixth phase of CRISP–DM focuses on the deployment of the proposed cash flow underwriting workflow in a production environment. This workflow is designed to leverage verifiable transaction data to enable transparent and evidence-based credit decisions while establishing a continuous feedback cycle for data collection, model retraining, and performance monitoring within existing core banking systems. To support this, a robust machine learning operations framework is implemented to structure the end-to-end pipeline into five key stages. Lastly, an integrated credit scoring framework is introduced to provide comprehensive risk evaluation for both established and new-to-credit MSMEs.
Data Ingestion
@@
The pipeline begins when bank statements are received. Extraction modules parse and structure the raw transaction data using OCR and layout analysis techniques. This process transforms unstructured documents into a standardised tabular format suitable for downstream processing. The ingestion stage also performs integrity verification and data validation checks to ensure document authenticity.
Feature Engineering and Feature Store
@@
Following ingestion, structured data are fed into the feature engineering module, which computes transaction-derived and behavioural features. These features are time-stamped, versioned, and stored in a centralised repository. The feature store provides a single source of truth for both training and inference, ensuring the same logic used to generate historical features is consistently applied to new applicants in real time. This design guarantees reproducibility, prevents feature skew, and supports traceability across model iterations.
Continuous Integration (CI)
@@
The CI pipeline automates model retraining and validation prior to integration into the production registry. It is triggered by two primary events: (1) a predefined schedule (e.g., annually) when new labelled data become available from the core banking system, and (2) alerts from the monitoring system indicating potential model or data drift. Upon activation, the pipeline retrieves the latest versioned feature set and ground truth labels, retrains the model, and evaluates its performance against the current production model using predefined and consistent evaluation metrics.
Model Registry and Continuous Deployment (CD)
@@
If the retrained model demonstrates superior performance, it is automatically versioned and stored in the model registry. The CD process then packages the validated model into a containerised microservice and deploys it as a secure REST API endpoint. To ensure reliability and minimise production risk, a canary deployment strategy is applied—initially routing a small fraction of live traffic to the new model for performance validation before full-scale rollout.
Continuous Monitoring and Feedback Loop
@@
Once the bank statement-based credit scoring model is deployed, it operates within a Champion–Challenger framework (kim2019champion) to enable continuous performance optimisation. The Champion model serves as the current production baseline, while one or more Challenger models are periodically retrained on the latest data and evaluated in parallel. Model performance is continuously monitored and compared across predefined metrics. When a Challenger consistently outperforms the Champion, the system automatically notifies designated bank officers to review the results and promote the new model to production. This approach ensures sustained model improvement, robustness, and operational stability to maintain optimal performance.
Integrated Credit Scoring Framework
@@
In production, the cash flow underwriting system functions as an integrated decision engine embedded within the lending process. Upon submission of bank statements, AI modules automatically extract and structure the raw data using OCR and natural language processing models. The processed data then undergoes fraud detection, cash flow analysis, and network analytics to generate transaction-derived features and credit scores for MSME applicants. The existing bureau-based scorecard, designed for borrowers with established credit histories, operates in parallel with the newly developed cash flow–based model. Each model independently produces a risk rating, and a risk override mechanism ensures conservative decisioning; if either model indicates higher risk, the final classification adopts that rating. This framework provides a transparent, evidence-based, and data-driven approach to MSME credit assessment using adaptive, explainable AI scoring. It also extends financial access to new-to-credit businesses through verifiable transaction data.
8 Conclusion
This study is the first to introduce bank statement transactions as alternative data for MSME loan underwriting in Malaysia. Our study demonstrates that transaction-derived features capture dynamic aspects of MSMEs’ financial behaviour that are often overlooked in traditional credit models. These features also exhibit strong predictive power with high IV scores. Empirical results show that models built solely on application information perform modestly, whereas those based on bank transaction data achieve substantially better results. The combined features further improve the model’s predictive power, confirming the complementary nature and practical value of alternative data.
9 Acknowledgments
We thank the Cradle Fund and the Microsoft for Startups program for their continuous support. This work was made possible through the collective contributions of the DocLab, DocSuite and MLSuite teams at AI Lens, whose efforts were instrumental in the development of this research.
10 Limitations and Future Work
@@
This work has several limitations that should be addressed in future research. First, the dataset comprises 611 loan applications from a single Malaysian lending institution. While the dataset represents real-world production data collected over a two-year period, the sample size is constrained by the practical realities of MSME lending in emerging markets, where data availability remains limited due to stringent data privacy and regulatory requirements. Future work should validate these findings across multiple institutions and larger datasets as additional MSME lending data become available.
@@
Second, the class imbalance in our dataset (518 non-default vs. 93 default cases) reflects the natural distribution observed in real-world lending portfolios, where the majority of borrowers successfully repay their loans. While this imbalance poses modelling challenges, it accurately reflects real-world credit risk settings. In this work, we addressed this limitation through appropriate model selection, WOE-based feature engineering, and careful evaluation using AUROC, which is robust to class imbalance. It is important to note that class imbalance is not a limitation that can be fully eliminated, but rather an inherent characteristic of credit scoring datasets. As credit assessment models improve and lending decisions become more accurate, successful repayment rates increase, thereby maintaining or even intensifying this natural imbalance. Nevertheless, future studies could explore advanced resampling techniques or cost-sensitive learning methods to further enhance model performance on minority class prediction.
@@
Third, while we describe multiple AI modules for document processing, fraud detection, and transaction analysis within the proposed end-to-end cash flow underwriting workflow, the evaluation focuses on overall credit scoring performance rather than the effectiveness of individual modules. However, detailed module-level assessment is constrained by the use of proprietary methods and sensitive operational data that cannot be publicly disclosed. Future work utilising synthetic or anonymised data could enable more granular module-level analysis and benchmarking. Additionally, future research could explore formalising these modules as fully autonomous agents, enabling the development of a multi-agent credit scoring architecture with clearer inter-agent responsibilities and interactions.
@@
Fourth, the proposed workflow has been deployed in a production environment with one Malaysian lending institution. While this demonstrates real-world applicability, broader validation across different lending institutions, regulatory environments, and MSME segments would strengthen the generalisability of our findings. Further research is needed to systematically integrate the proposed cash flow underwriting workflow with existing credit assessment systems that rely on behavioural and credit bureau data, ensuring coherent and robust risk evaluation. Additionally, longitudinal studies examining model performance stability across economic cycles would provide valuable insights into the temporal robustness and adaptability of bank statement-based features under varying macroeconomic conditions.

### v4

@@ Appendix B Deployment and Operational Framework
The sixth phase of CRISP-DM focuses on the deployment of the proposed cash flow underwriting workflow in a production environment. This workflow is designed to leverage verifiable transaction data to enable transparent and evidence-based credit decisions while establishing a continuous feedback cycle for data collection, model retraining, and performance monitoring within existing core banking systems. To support this, a robust machine learning operations framework is implemented to structure the end-to-end pipeline into five key stages. Lastly, an integrated credit scoring framework is introduced to provide comprehensive risk evaluation for both established and new-to-credit MSMEs.
B.1 Data Ingestion
@@
The pipeline begins when bank statements are received. Extraction modules parse and structure the raw transaction data using OCR and layout analysis techniques. This process transforms unstructured documents into a standardized tabular format suitable for downstream processing. The ingestion stage also performs integrity verification and data validation checks to ensure document authenticity.
B.2 Feature Engineering and Feature Store
@@
Following ingestion, structured data are fed into the feature engineering module, which computes transaction-derived and behavioral features. These features are time-stamped, versioned, and stored in a centralized repository. The feature store provides a single source of truth for both training and inference, ensuring the same logic used to generate historical features is consistently applied to new applicants in real time. This design guarantees reproducibility, prevents feature skew, and supports traceability across model iterations.
B.3 Continuous Integration (CI)
@@
The CI pipeline automates model retraining and validation prior to integration into the production registry. It is triggered by two primary events: (1) a predefined schedule (e.g., annually) when new labeled data become available from the core banking system, and (2) alerts from the monitoring system indicating potential model or data drift. Upon activation, the pipeline retrieves the latest versioned feature set and ground truth labels, retrains the model, and evaluates its performance against the current production model using predefined and consistent evaluation metrics.
B.4 Model Registry and Continuous Deployment (CD)
@@
If the retrained model demonstrates superior performance, it is automatically versioned and stored in the model registry. The CD process then packages the validated model into a containerized microservice and deploys it as a secure REST API endpoint. To ensure reliability and minimize production risk, a canary deployment strategy is applied, initially routing a small fraction of live traffic to the new model for performance validation before full-scale rollout.
B.5 Continuous Monitoring and Feedback Loop
@@
Once the bank statement-based credit scoring model is deployed, it operates within a Champion-Challenger framework (Kim et al., 2019) to enable continuous performance optimization. The Champion model serves as the current production baseline, while one or more Challenger models are periodically retrained on the latest data and evaluated in parallel. Model performance is continuously monitored and compared across predefined metrics. When a Challenger consistently outperforms the Champion, the system automatically notifies designated bank officers to review the results and promote the new model to production. This approach ensures sustained model improvement, robustness, and operational stability.
B.6 Integrated Credit Scoring Framework
@@
In production, the cash flow underwriting system functions as an integrated decision engine embedded within the lending process. Upon submission of bank statements, AI modules automatically extract and structure the raw data using OCR and natural language processing models. The processed data then undergoes fraud detection, cash flow analysis, and network analytics to generate transaction-derived features and credit scores for MSME applicants. The existing bureau-based scorecard, designed for borrowers with established credit histories, operates in parallel with the newly developed cash flow-based model. Each model independently produces a risk rating, and a risk override mechanism ensures conservative decisioning; if either model indicates higher risk, the final classification adopts that rating. This framework provides a transparent, evidence-based, and data-driven approach to MSME credit assessment using adaptive, explainable AI scoring. It also extends financial access to new-to-credit businesses through verifiable transaction data.
Appendix C Limitations and Future Work
@@
This work has several limitations that should be addressed in future research. First, the dataset comprises 611 loan applications from a single Malaysian consulting firm. While the dataset represents real-world production data collected over a two-year period, the sample size is constrained by the practical realities of MSME lending in emerging markets, where data availability remains limited due to stringent data privacy and regulatory requirements. Future work should validate these findings across multiple institutions and larger datasets as additional MSME lending data become available.
@@
Second, the class imbalance in our dataset (518 non-default vs. 93 default cases) reflects the natural distribution observed in real-world lending portfolios, where the majority of borrowers successfully repay their loans. While this imbalance poses modeling challenges, it accurately reflects real-world credit risk settings. In this work, we addressed this limitation through appropriate model selection, WOE-based feature engineering, and careful evaluation using AUROC, which is robust to class imbalance. It is important to note that class imbalance is not a limitation that can be fully eliminated, but rather an inherent characteristic of credit scoring datasets. As credit assessment models improve and lending decisions become more accurate, successful repayment rates increase, thereby maintaining or even intensifying this natural imbalance. Nevertheless, future studies could explore advanced resampling techniques or cost-sensitive learning methods to further enhance model performance on minority class prediction.
@@
Third, while we describe multiple AI modules for document processing, fraud detection, and transaction analysis within the proposed end-to-end cash flow underwriting workflow, the evaluation focuses on overall credit scoring performance rather than the effectiveness of individual modules. However, detailed module-level assessment is constrained by the use of proprietary methods and sensitive operational data that cannot be publicly disclosed. Future work utilizing synthetic or anonymized data could enable more granular module-level analysis and benchmarking. Additionally, future research could explore formalizing these modules as fully autonomous agents, enabling the development of a multi-agent credit scoring architecture with clearer inter-agent responsibilities and interactions.
@@
Fourth, the proposed workflow has been deployed in a production environment with one Malaysian consulting firm. While this demonstrates real-world applicability, broader validation across different lending institutions, regulatory environments, and MSME segments would strengthen the generalizability of our findings. Further research is needed to systematically integrate the proposed cash flow underwriting workflow with existing credit assessment systems that rely on behavioral and credit bureau data, ensuring coherent and robust risk evaluation. Additionally, longitudinal studies examining model performance stability across economic cycles would provide valuable insights into the temporal robustness and adaptability of bank statement-based features under varying macroeconomic conditions.